# Bayesian Optimization with Formal Safety Guarantees via Online Conformal Prediction

### v1

@@ I-C Main Contributions
TABLE I: State of the art on Safe-BO against the proposed Safe-BOCP
Safe-BO [7, 10, 8, 9, 11] SAFE-BOCP (ours)
Target safety violation rate 0 (0,1]
Assumption-free safety guarantee
@@ II-A Optimization Problem and Safety Constraint
We focus on optimization problems constrained of the form
@@
$\bullet$ Safety: Interpreting the inequality $q(\mathbf{x})\geq 0$ as a safety constraint, we consider choices of the optimization variable $\mathbf{x}\in\mathcal{X}$ that result in a negative value of the constraint function $q(\mathbf{x})$ to be unsafe, unless the number of such violations of the constraint are kept below a threshold. Accordingly, we will require that the number of evaluations of the constraint function $q(\mathbf{x})$ that result in a violation of the constraint in (1) be no larger than a pre-determined value. We will formalize this constraint next by describing the general operation of the optimizer.
@@ II-B Sequential Surrogate-Based Safe Optimization
where the observation noise terms $\epsilon_{f,t}\sim\mathcal{N}(0,\sigma_{f}^{2})$ and $\epsilon_{q,t}\sim\mathcal{N}(0,\sigma_{q}^{2})$ are independent zero-mean Gaussian random variables with variance $\sigma_{f}^{2}$ and $\sigma_{q}^{2}$, respectively [7, 10, 8, 9, 11].
@@ II-C2 Probabilistic Safety Constraint
In the presence of observation noise on the constraint, i.e., with a positive observation noise power $\sigma_{q}^{2}>0$, the optimizer cannot guarantee the deterministic constraint (7). Rather, for some target violation rate $\alpha\in[0,1]$, the optimizer can only aim at ensuring that the constraint (7) be satisfied with a probability no smaller than a target reliability level $1-\delta$, with $\delta\in(0,1]$. This results in the probabilistic safety constraint
@@ IV-B Safe Set Creation
By choosing the scaling parameter $\beta_{t+1}$ as in (22), under the key assumption (21), all the decisions in the safe set $\mathcal{S}_{t+1}$ (17) can be proved to be safe with high probability [8, Lemma 1] (see also [25, Theorem 6]).
@@ IV-C Acquisition Process
$\mathcal{M}_{t+1}=\bigg{\{}\mathbf{x}\in\mathcal{S}_{t+1}\Big{|}f_{u}(\mathbf{x}|\mathcal{O}_{t})\geq\max\limits_{\mathbf{x}^{\prime}\in S_{t+1}}f_{l}(\mathbf{x}|\mathcal{O}_{t})\bigg{\}}.$ (25)
@@
and the set of possible expanders is defined as
The overall procedure of SAFEOPT is summarized in Algorithm 1.
@@ V Deterministic Safe-BO via Online Conformal Prediction
As we have reviewed in Sec. IV, in order to achieve a zero target violation rate $\alpha=0$ in the safety constraints (7) and (8), SAFEOPT assumes that the constraint function $q(\mathbf{x})$ belongs to a specific family of functions. Other Safe-BO methods [10, 8, 9] also require the same assumption to guarantee the safety constraint (see Sec. I). In the following two sections, we will introduce SAFE-BOCP, a novel Safe-BO scheme that achieves the safety constraint requirements (7) or (8) without requiring any assumptions on the underlying constraint function $q(\mathbf{x})$. This goal is met at the cost of obtaining a non-zero, controllable, target violation rate $\alpha\in(0,1]$ in the deterministic safety requirement (7) and in the probabilistic safety requirement (8). This section focuses on the case in which observations (5) of the constraint function are ideal, i.e., $\sigma_{q}^{2}=0$, hence aiming at achieving the deterministic safety constraint (7). The next section addresses the case with noisy observations on the constraint function.
@@ V-C Safety Guarantees
       Evaluate error signal $\text{err}_{t}=\mathbbm{1}(z_{t}<0)$ as in (29)
@@
Algorithm 2 D-SAFE-BOCP
VI Probabilistic Safe-BOCP
We now turn to the case in which the observations (5) of constraint function $q(\mathbf{x})$ are noisy ($\sigma_{q}^{2}>0$). The main challenge in extending the approach proposed in the previous section is the fact that the error signal (29) is an unreliable indication of whether candidate $\mathbf{x}_{t}$ is safe or not due to the presence of observation noise. Accordingly, we start by proposing an alternative way to measure the excess violation rate.
VI-A P-SAFE-BOCP
The main idea underlying the proposed P-SAFE-BOCP is to count as unsafe all solutions $\mathbf{x}_{t}$ for which the noisy observation $z_{t}=q(\mathbf{x}_{t})+\epsilon_{q,t}$ in (5) is smaller than some positive threshold $\sigma_{q}\cdot\omega$ that scales with the observation noise standard deviation $\sigma_{q}$ with proportionality constant $\omega>0$. Accordingly, the safety error signal is defined as
$\text{err}_{t}=\mathbbm{1}(z_{t}<\sigma_{q}\cdot\omega).$ (39)
In a manner consistent with the previous section, for noiseless observations, we recover the error signal (29) used by D-SAFE-BOCP, since we have a zero observation noise power $\sigma^{2}_{q}=0$.
More generally, when the observation noise power $\sigma^{2}_{q}$ is positive, the threshold $\sigma_{q}\cdot\omega$ used in the error signal (39) causes P-SAFE-BOCP to cautiously treat as errors all observations $z_{t}$ that are not sufficiently large to overcome the uncertainty associated with the noisy observation of the constraint function.
Intuitively, the slope $\omega$ should increase with the target reliability level $1-\delta$ in the probabilistic safety constraint (8). In fact, a larger target reliability level calls for more caution in determining whether a given observation $z_{t}$ of the constraint function is likely to indicate an unsafe solution or not. In line with this intuition, we set the slope factor $\omega$ as
$\omega=F^{-1}\big{(}(1-\delta)^{\frac{1}{T}}\big{)},$ (40)
where $F^{-1}(\cdot)$ is the inverse CDF of standard Gaussian distribution (34).
The reason for this choice is elucidated by the following lemma, which relates the true violation rate (7) to the estimated violation rate $\sum_{t=1}^{T}\text{err}_{t}/T$ using the error signal (39).
@@ VI-A P-SAFE-BOCP
[…]
@@
When a candidate solution $\mathbf{x}_{t}$ is unsafe, i.e., when $q(\mathbf{x}_{t})<0$, the probability that the error signal $\text{err}_{t}$ in (39) correctly reports an error, setting $\text{err}_{t}=1$, is lower bonded by $F(\omega)$, where $F(\cdot)$ is the CDF of standard Gaussian distribution. Therefore, the probability that the true violation rate violation-rate($T$) no larger than the estimated violation rate $\sum_{t=1}^{T}\text{err}_{t}/T=1$ is lower bounded by the probability that all the errors correctly reported. This is, in turn, lower bounded by $(F(\omega))^{T}$ by the independence of the observation noise variables $\{\epsilon_{q,t}\}_{t=1}^{T}$.
As specified in Algorithm 3, P-SAFE-BOCP follows the same steps in D-SAFE-BOCP with the caveat that the error signal (39) is used in lieu of (29). As we prove next, the correction applied via the error signal (39) is sufficient to meet the probabilistic safety requirement (8).
@@ VI-B Safety Guarantees
The satefy guarantees of P-SAFE-BOCP are summarized in the following theorem.
@@
Under noisy observations of the constraint function ($\sigma_{q}^{2}>0$), P-SAFE-BOCP satisfies the probabilistic safety constraint (8) for any pre-determined target violation rate $\alpha\in(0,1]$ and target reliability level $\delta\in(0,1)$.
@@
Input : GP priors $(\mu_{f}(\mathbf{x}),\kappa_{f}(\mathbf{x},\mathbf{x}^{\prime}))$ and $(\mu_{q}(\mathbf{x}),\kappa_{q}(\mathbf{x},\mathbf{x}^{\prime}))$, initial safe set $\mathcal{S}_{0}$, initial observation $\mathcal{O}_{0}$, total number of optimization iterations $T$, target violation rate $\alpha$, update rate $\eta>0$, initial excess violation rate $\Delta\alpha_{1}<1$, observation noise power $\sigma_{q}^{2}$ for the constraint
Output : Decision $\mathbf{x}^{*}$
 Initialize $\mathbf{x}_{1}=\textsf{S{\footnotesize AFE}O{\footnotesize PT}}(\mathcal{O}_{0}|\beta_{1})$ using $\beta_{1}=\varphi(\Delta\alpha_{1})$ (34), algorithmic target level $\alpha_{\text{algo}}$ as in (35), slope factor $\omega$ as in (40)
for $t=1,...T$ do
       Observe $y_{t}$ and $z_{t}$ from candidate solution $\mathbf{x}_{t}$
       Update GPs with $\mathcal{O}_{t}$ as in (12) and (13)
       Evaluate cautious error signal $\text{err}_{t}=\mathbbm{1}(z_{t}<\sigma_{q}\cdot\omega)$ as in (39)
       $\mathbf{x}_{t+1}=$SAFEOPT$(\mathcal{O}_{t}|\beta_{t+1})$ from Algorithm 1
end for
Algorithm 3 P-SAFE-BOCP
@@ VII-A Synthetic Dataset
As discussed throughout the paper, the scaling parameter for the constraint function $q(\mathbf{x})$ in (15) is a priori determined by (22) for SAFEOPT, and is adapted by feedback via $\beta_{t+1}=\varphi(\Delta\alpha_{t+1})$ (30) for the proposed SAFE-BOCP, while we fix the scaling parameter for the objective function $f(\mathbf{x})$ in (14) to $3$ since it does not affect the safety guarantee for both SAFEOPT (see [25, Theorem 6]) and SAFE-BOCP. The objective observation noise variance is set to $\sigma_{f}^{2}=2.5\times 10^{-3}$; the initial safe decision is chosen as $\mathbf{x}_{0}=0$ for which we have $q(\mathbf{x}_{0})=0.946>0$; and the total number of optimization iterations is set to $T=25$. For SAFE-BOCP, we set the update rate in (30) to $\eta=2.0$. All results are averaged over 1,000 experiments, with error bars shown to encompass $95\%$ of the realizations. Each experiment corresponds to a random draw of the objective function and to random realization of the observation noise signals.
@@ VII-B Deterministic Safety Requirement
As explained in Sec. IV, SAFEOPT requires the GP model for the constraint function $q(\mathbf{x})$ to be well specified (18)–(21) in order to meet safety conditions. To study the impact of violations of this assumption, we start by considering the noiseless case, i.e., $\sigma_{q}^{2}=0$, and we vary the kernel bandwidth $h$ adopted for the GP models used as surrogates for the objective and constraint functions as discussed earlier. In this experiment, we set the target violation rate in the deterministic safety constraint (7) as $\alpha=0.1$ for D-SAFE-BOCP, while we recall that SAFEOPT assumes the target $\alpha=0$.
@@
In terms of optimality ratio, SAFEOPT in the “safe” regime $B/||q||_{\kappa}\geq 1$ with well-specified GP parameter $h$, achieves around 85%, while D-SAFE-BOCP obtains 84.5%. With a misspecified value $h$, D-SAFE-BOCP achieves an optimality ratio around 87.5%, while the optimality ratio of SAFEOPT is larger, but not relevant given the violation of the safety requirement. Note that a misspecified value of the kernel bandwidth $h$ does not necessarily reduce the performance of D-SAFE-BOCP, which is improved in this example.
[…]
@@
The figure shows that the violation rate obtained by SAFEOPT is above its target $\alpha=0$. In contrast, as per the theory developed in this paper, D-SAFE-BOCP obtains flexible and guaranteed violation rates within the tolerated ranges for all values of $\alpha$. Furthermore, as the tolerated violation rate $\alpha$ increases, the optimality ratio of D-SAFE-BOCP is enhanced, indicating a trade-off between the two metrics.

### v3

@@ I-B Related Work
In a related line of work, the constrained BO approaches in [21, 22, 23] target a constrained optimization problem, but allow unlimited safety violations during the optimization process. More recent references [24, 25] considered an explicit budget of safety violations for probabilistic constraints, but did not provide any formal safety guarantees.
@@
TABLE I: State of the art on Safe-BO against the proposed Safe-BOCP
Safe-BO [7, 10, 8, 9, 11] SAFE-BOCP (ours)
Target safety violation rate 0 (0,1]
Assumption-free safety guarantee
@@ II-A Optimization Problem and Safety Constraint
We focus on constrained optimization problems of the form
@@
$\bullet$ Safety: Interpreting the inequality $q(\mathbf{x})\geq 0$ as a safety constraint, we consider choices of the optimization variable $\mathbf{x}\in\mathcal{X}$ that result in a negative value of the constraint function $q(\mathbf{x})$ to be unsafe, unless the number of such violations of the constraint are kept below a threshold. Accordingly, we will require that the number of evaluations of the constraint function $q(\mathbf{x})$ that result in a violation of the inequality $q(\mathbf{x})\geq 0$ to be no larger than a pre-determined value. We will formalize this constraint next by describing the general operation of the optimizer.
@@ II-B Sequential Surrogate-Based Safe Optimization
where the observation noise for the objective, $\epsilon_{f,t}\sim\mathcal{N}(0,\sigma_{f}^{2})$, is Gaussian with variance $\sigma_{f}^{2}$, while the observation noise for the constraint, $\epsilon_{q,t}$, can follow any distribution provided that it has a known upper bound on the one-sided right-tail probability (see Assumption 1 in Sec. VI-A for details).
@@ II-C2 Probabilistic Safety Constraint
In the presence of observation noise on the constraint, i.e., with a positive observation noise power $\sigma_{q}^{2}>0$, the optimizer cannot guarantee the deterministic constraint (7). Rather, targeting problem (1), the optimizer can only aim at ensuring that the constraint (7) be satisfied with a probability no smaller than a target reliability level $1-\delta$, with $\delta\in(0,1]$. This results in the probabilistic safety constraint
@@
in which the probability is taken with respect to the observation noise variables $\{\epsilon_{q,t}\}_{t=1}^{T}$ for the constraint function $q(\mathbf{x})$ in (5). Therefore, in this second case, we target the maximization of function $f(x)$ subject to the safety constraint (8) on the optimization process.
@@ IV-B Safe Set Creation
By choosing the scaling parameter $\beta_{t+1}$ as in (22), under the key assumption (21), all the decisions in the safe set $\mathcal{S}_{t+1}$ (17) can be proved to be safe with high probability [8, Lemma 1] (see also [30, Theorem 6]).
@@ IV-C Acquisition Process
$\mathcal{M}_{t+1}=\bigg{\{}\mathbf{x}\in\mathcal{S}_{t+1}\Big{|}f_{u}(\mathbf{x}|\mathcal{O}_{t})\geq\max\limits_{\mathbf{x}^{\prime}\in S_{t+1}}f_{l}(\mathbf{x}^{\prime}|\mathcal{O}_{t})\bigg{\}}.$ (25)
@@
and the set of possible expanders is defined as
The overall procedure of SAFEOPT is summarized in Algorithm 1.
@@ V Deterministic Safe-BO via Online Conformal Prediction
As we have reviewed in Sec. IV, in order to achieve a zero target violation rate $\alpha=0$ in the safety constraints (7) and (8), SAFEOPT assumes that the constraint function $q(\mathbf{x})$ belongs to a specific family of functions. Other Safe-BO methods [10, 8, 9] also require the same assumption to guarantee the safety constraint (see Sec. I). In the following two sections, we will introduce SAFE-BOCP, a novel Safe-BO scheme that achieves the safety constraint requirements (7) or (8) without requiring any assumptions on the underlying constraint function $q(\mathbf{x})$. This goal is met at the cost of obtaining a non-zero, controllable, target violation rate $\alpha\in(0,1]$ in the deterministic safety requirement (7) and in the probabilistic safety requirement (8). This section focuses on the case in which observations (5) of the constraint function are ideal, i.e., $\epsilon_{q,t}=0$, hence aiming at achieving the deterministic safety constraint (7). The next section addresses the case with noisy observations on the constraint function.
@@ V-B D-SAFE-BOCP
Output : Decision $\mathbf{x}^{*}$
for $t=1,...T$ do
       Observe $y_{t}$ and $z_{t}$ from candidate solution $\mathbf{x}_{t}$
       Update GPs with $\mathcal{O}_{t}$ as in (12) and (13)
       Evaluate error signal $\text{err}_{t}=\mathbbm{1}(z_{t}<0)$ as in (29)
       $\mathbf{x}_{t+1}=$SAFEOPT$(\mathcal{O}_{t}|\beta_{t+1})$ from Algorithm 1
end for
Algorithm 2 D-SAFE-BOCP
@@ V-C Safety Guarantees
VI Probabilistic Safe-BOCP
We now turn to the case in which the observations (5) of constraint function $q(\mathbf{x})$ are noisy. The main challenge in extending the approach proposed in the previous section is the fact that the error signal (29) is an unreliable indication of whether candidate $\mathbf{x}_{t}$ is safe or not due to the presence of the observation noise $\epsilon_{q,t}$. Accordingly, we start by proposing an alternative way to measure the excess violation rate.
VI-A P-SAFE-BOCP
To proceed, we assume that the observation noise $\epsilon_{q,t}$ in (5) has a known upper bound on the right-tail probability $\text{Pr}(\epsilon_{q,t}\geq\omega)$ for all $\omega\in\mathbbm{R}$. This basic assumption is also adopted in the robust CP literature [31, Theorem 1]. In Sec. VI-C, we will illustrate how to further alleviate this assumption by assuming access to noise samples.
Assumption 1.
The constraint observation noise $\epsilon_{q,t}$, which is independent over $t=1,...,T$, has a known upper bound $F^{+}(\omega)$ on its one-sided right-tail probability, i.e.,
$\text{Pr}(\epsilon_{q,t}\geq\omega)\leq F^{+}(\omega)$ (39)
for all $t=1,...,T$ and any $\omega\in\mathbb{R}$.
The main idea underlying the proposed P-SAFE-BOCP is to count as unsafe all solutions $\mathbf{x}_{t}$ for which the noisy observation $z_{t}=q(\mathbf{x}_{t})+\epsilon_{q,t}$ in (5) is smaller than some back-off threshold $\omega_{q}>0$. Specifically, we define the safety error signal as
$\text{err}_{t}=\mathbbm{1}(z_{t}<\omega_{q}),$ (40)
where the corresponding threshold $\omega_{q}$ is obtained as
$\omega_{q}=\inf\{\omega\in\mathbb{R}:F^{+}(\omega)\leq 1-(1-\delta)^{\frac{1}{T}}\}.$ (41)
The threshold $\omega_{q}$ increases with the target reliability level $1-\delta$ in the probabilistic safety constraint (8). In fact, a larger target reliability level calls for more caution in determining whether a given observation $z_{t}$ of the constraint function is likely to indicate an unsafe solution or not.
The rationale behind the definitions (40)-(41) is formalized by the following lemma, which relates the true violation rate (7) to the estimated violation rate $\sum_{t=1}^{T}\text{err}_{t}/T$ using the error signal (40).
@@ VI-A P-SAFE-BOCP
       Evaluate cautious error signal $\text{err}_{t}=\mathbbm{1}(z_{t}<\omega_{q})$ as in (40) with $\omega_{q}$ obtained from (41)
@@
Algorithm 3 P-SAFE-BOCP
@@
$\Pr\bigg{(}\text{violation-rate($T$)}\leq\frac{1}{T}\sum_{t=1}^{T}\text{err}_{t}\bigg{)}$ $\geq(1-F^{+}(\omega))^{T}$
$=1-\delta,$ (42)
@@
When a candidate solution $\mathbf{x}_{t}$ is unsafe, i.e., when $q(\mathbf{x}_{t})<0$, the probability that the error signal $\text{err}_{t}$ in (40) correctly reports an error, setting $\text{err}_{t}=1$, is lower bounded by $1-F^{+}(\omega)$. Therefore, the probability that the true violation rate violation-rate($T$) no larger than the estimated violation rate $\sum_{t=1}^{T}\text{err}_{t}/T=1$ is lower bounded by the probability that all the errors correctly reported. This is, in turn, lower bounded by $(1-F^{+}(\omega))^{T}$ by the independence of the observation noise variables $\{\epsilon_{q,t}\}_{t=1}^{T}$. ∎
As specified in Algorithm 3, P-SAFE-BOCP follows the same steps in D-SAFE-BOCP with the caveat that the error signal (40) is used in lieu of (29). As we will prove next, the correction applied via the safety error signal (40) is sufficient to meet the probabilistic safety requirement (8).
@@ VI-B Safety Guarantees
The safety guarantee of P-SAFE-BOCP is summarized in the following theorem.
@@
Under noisy observations of the constraint function and Assumption 1, P-SAFE-BOCP satisfies the probabilistic safety constraint (8) for any pre-determined target violation rate $\alpha\in(0,1]$ and target reliability level $\delta\in(0,1)$.
@@
VI-C Data-Driven Probability Bound
A possible challenge in applying P-SAFE-BOCP in practice is the fact that an upper bound $F^{+}(\omega)$ on the probability $\text{Pr}(\epsilon_{q,t}\geq\omega)$ may not be known a priori. In this subsection, we provide a data-driven approach for evaluating an upper bound on the probability $\text{Pr}(\epsilon_{q,t}\geq\omega)$, assuming only access to independent and identically distributed (i.i.d.) observation noise samples.
Lemma 2 (Estimated Upper Bound).
Assume access to i.i.d. observation noise samples $\{\epsilon_{q,i}\}_{i=1}^{m}$. The empirical estimate of the right-tail probability
$\hat{F}^{+}(\omega)=\frac{1}{m}\sum_{i=1}^{m}\mathbbm{1}(\epsilon_{q,i}>\omega),$ (45)
when offset by $\psi>0$, provides an upper bound on $\text{Pr}(\epsilon_{q,i}>\omega)$ with probability
$\Pr\big{(}\Pr(\epsilon_{q,i}\geq\omega,\forall\omega\in\mathbbm{R})\leq\hat{F}^{+}(\omega)+\psi\big{)}$
$\geq 1-\exp(-2m\psi^{2})$ (46)
for any $\psi>\sqrt{\ln 2/2m}$.
Lemma 2 is a direct application of Dvoretsky-Kiefer-Wolfowitz inequality [32].
Consequently, by using $\hat{F}^{+}(\omega)+\psi$ in lieu of $F^{+}(\omega)$ in (41), we have the following modified safety guarantee of P-SAFE-BOCP.
Corollary 1.
Under noisy observations of the constraint function, P-SAFE-BOCP with $\hat{F}^{+}(\omega)+\psi$, for any $\psi>0$, in lieu of $F^{+}(\omega)$ in (41) satisfies the guarantee
$\Pr(\text{violation-rate}(T)\leq\alpha)\geq(1-\exp{(-2m\psi^{2})})(1-\delta)$ (47)
for any pre-determined target violation rate $\alpha\in(0,1]$ and target reliability level $\delta\in(0,1)$, where the probability is taken with respect to the observation noise variables $\{\epsilon_{q,t}\}_{t=1}^{T}$ as well as the $m$ i.i.d. noise samples in (45).
Corollary 1 is obtained by combining Lemma 2 and Theorem 3. Intuitively, with an increasing number $m$ of the constraint observation noise samples, the tightness of the safety guarantee in Theorem 3 is enhanced as a result of increasingly accurate observation noise estimation.
@@ VII-A Synthetic Dataset
Figure 4: Violation-rate$(t)$ (top) and optimality ratio (bottom) against iteration index $t$ with target violation rate $\alpha=0.3$ (dot-dashed line), update rate $\eta=2$, misspecified kernel bandwidth $h=1/14.58$, RKHS norm bound $B=||q||_{\kappa^{*}}$ and total number of iteration $T=50$.
As discussed throughout the paper, the scaling parameter for the constraint function $q(\mathbf{x})$ in (15) is a priori determined by (22) for SAFEOPT, and is adapted by feedback via $\beta_{t+1}=\varphi(\Delta\alpha_{t+1})$ (30) for the proposed SAFE-BOCP, while we fix the scaling parameter for the objective function $f(\mathbf{x})$ in (14) to $3$ since it does not affect the safety guarantee for both SAFEOPT (see [30, Theorem 6]) and SAFE-BOCP. The objective observation noise variance is set to $\sigma_{f}^{2}=2.5\times 10^{-3}$; and the initial safe decision is chosen as $\mathbf{x}_{0}=0$ for which we have $q(\mathbf{x}_{0})=0.946>0$. For SAFE-BOCP, we set the update rate in (30) to $\eta=2.0$. All results are averaged over 1,000 experiments, with error bars shown to encompass $95\%$ of the realizations. Each experiment corresponds to a random draw of the objective function and to random realization of the observation noise signals. The implementation examples of the synthetic benchmark can be found here [^1] https://github.com/yunchuan-zhang/safe-bocp.
@@ VII-B Deterministic Safety Requirement
As explained in Sec. IV, SAFEOPT requires the GP model for the constraint function $q(\mathbf{x})$ to be well specified (18)–(21) in order to meet safety conditions. To study the impact of violations of this assumption, we start by considering the noiseless case, i.e., $\sigma_{q}^{2}=0$, and we vary the kernel bandwidth $h$ adopted for the GP models used as surrogates for the objective and constraint functions as discussed earlier.
Fig. 4 shows the violation rate and optimality ratio against the iteration index $t$. For D-SAFE-BOCP, we set the update rate as $\eta=2$ and the target violation rate to $\alpha=0.3$, while SAFEOPT assumes target $\alpha=0$ with RKHS norm bound $B=||q||_{\kappa^{*}}$. For both schemes, the total number of iterations is $T=50$, and the misspecified GP with RBF kernel bandwidth $h=1/14.58$ is adopted.
The violation rate obtained by SAFEOPT is above the target $\alpha=0.3$ for a significant interval of time $t$, and it progressively falls below the target with a larger $t$, while D-SAFE-BOCP meets the deterministic safety requirement (7) with the pre-determined target $\alpha=0.3$ across all iterations. Furthermore, the optimality ratio obtained by D-SAFE-BOCP is larger than SAFEOPT after iteration $t=13$, converging to $97.5\%$ at iteration $t=20$. In contrast, SAFEOPT converges to optimality ratio of $94.5\%$ at iteration $t=25$, at which point the target safety level $\alpha=0.3$ is violated.
[…]
@@
In terms of optimality ratio, in the regime $B/||q||_{\kappa^{*}}\geq 1$, with a well-specified GP parameter $h$, SAFEOPT achieves around 83%, while D-SAFE-BOCP obtains the larger optimality ratio 84.5%. In contrast, with a misspecified value $h$, D-SAFE-BOCP achieves an optimality ratio around 87.5%, while the optimality ratio of SAFEOPT is larger, but this comes at the cost of the violation of the safety requirement. Note that a misspecified value of the kernel bandwidth $h$ does not necessarily reduce the performance of D-SAFE-BOCP, which is improved in this example.
@@
The figure shows that the violation rate obtained by SAFEOPT exceeds its target $\alpha=0$, and thus the safety requirement is violated. In contrast, as per the theory developed in this paper, D-SAFE-BOCP meets violation-rate requirement for all values of the target $\alpha$. Moreover, as the tolerated violation rate $\alpha$ increases, the optimality ratio of D-SAFE-BOCP is enhanced, indicating a trade-off between the two metrics. When increasing the target violation rate $\alpha$, D-SAFE-BOCP raises the algorithmic target level $\alpha_{\text{algo}}$ in (35), making it possible for the optimizer to reduce the time spent under the maximally cautious scaling $\beta_{t}=\infty$ in (33). Consequently, with a larger $\alpha$, the optimality ratio of D-SAFE-BOCP gains from more explorations of the objective function $f(\mathbf{x})$.